i NEED HELP TRYING TI FIGURE THIS OUT.
PERIMETER OF A RECTANGLE - THE WIDTH OF A RECTANGULAR PLAYGROUND IS 2X - 5 FEET, AND THE LENGHT IS 3X + 9 FEET. WRITE A POLYNOMIAL P(X) THAT REPRESENTS THE PERIMETER AND THEN EVALUATE THIS PERIMETER POLYNOMIAL IF X IS 4 FEET

Answer by ankor@dixie-net.com(22740)

THE WIDTH OF A RECTANGULAR PLAYGROUND IS 2X - 5 FEET, AND THE LENGHT IS 3X + 9 FEET. WRITE A POLYNOMIAL P(X) THAT REPRESENTS THE PERIMETER AND THEN EVALUATE THIS PERIMETER POLYNOMIAL IF X IS 4 FEET
:
Width = (2x-5)
:
Length = (3x+9)
:
P = 2L + 2W
:
P(x) = 2(3x+9) + 2(2x-5)
:
Multiply what's inside the brackets
P(x) = (6x + 18) + (4x-10)
:
P(x) = 6x + 4x + 18 - 10; combine like terms
:
P(x) = 10x + 8
:
If x = 4 ft, substitute for 4 for x
:
P(x) = 10(4) + 8
:
P(x) = 40 + 8
:
P(x) = 48 ft
:
:
Did this help?